2.0 moles of CaCl2 in particles


1
Expert's answer
2021-11-09T10:24:24-0500

Solution:

One mole of any substance contains 6.02×1023 particles.

Hence,

(2.0 mol) × (6.02×1023 particles / 1 mol) = 1.2×1024 particles

Number of particles = 1.2×1024 particles


Answer: 1.2×1024 particles